First Store prepares budgets quarterly. The following information is available for use in planning the second quarter budgetsCash dividends are declared during the third month of each quarter and are paid during the first month of the following quartRequired: Prepare and answer the following budgets and questions on the Part1Budget Solution Worksheet. Make sure you use cel4. Prepare a cash budget for each month of the second quarter ending June 30, 2020. Include budgeted borrowings and repaymentBudget #1 First Store Purchase Budget For the quarter ending June 30, 2020 Quarter Total April May June 50,000 Budgeted SalesBudget #3 First Store Schedule of Cash Disbursements For the quarter ending June 30, 2020 Quarter Total April May June PurchaBudget #5 First Store Budgeted Income Statement For the quarter ending June 30, 2020 Quarter June Total April May Sales Cost

Please show excel formula references/formulas
Answer #1

Answer is given below

for excel formula screen shot is given

For Q.No. 7 if interest rate is changed to 10% profit and loss account and balance sheet will change.

New budgeted interest expense will be $417 for the quarter

you can change the interest rate in below worksheet and check it

Sales Budgeted COGS -49% Desired ending inventory Total Needs Less Opening Stock Budgeted Purchase Purchase Budget April $50,

А в First Store Cash Budget For the Three months Ended June 30, 2019 April May 2000 =D60 =D34 =E34 =SUM(D49:050) =SUM(E49:E50

ТА В G H 67 68 First Store Budgeted Income Statement For the Three Months Ending June 30, 2020 April May June =D4 =D73*49% =D

G H Assets Current Assets Cash =F60 Accounts receivable =D116 Inventory =F20 Prepaid Insurance 500 Fixed Assets Property and